OpenStudy (anonymous):

Multiply. State any excluded values. q+5/-3 * -3q/q+2

OpenStudy (anonymous):

A. -3q^2 - 15q / 3 ; where q = 3 B. -3q - 15q^2 / 2q - 6 ; where q =3 C. q^2 - 15q / -3q ; where q = 0 D. -3q^2 - 15q / -3q - 6 ; where q = -2

OpenStudy (anonymous):

-3\[-3q ^{2}-15q/-3q-6\]

OpenStudy (anonymous):

excluded q=-2

OpenStudy (anonymous):

d